Insert two geometric means between 320 and 135.
Solve using Geometric Sequence

Answer by Alan3354(69443)

Insert two geometric means between 320 and 135.
----
320*g*g*g = 135
g^3 = 135/320 = 27/64
g = 3/4
------
--> 320, 240, 180, 135